OpenStudy (anonymous):

Find the polynomial function with roots 11 and 2i.

OpenStudy (anonymous):

(x-11)(x-2i)

OpenStudy (anonymous):

If complex allowed... Else assume -2i as well.

OpenStudy (anonymous):

ok but how do you do it

OpenStudy (anonymous):

Romero gave a complex polynomial, if you want real then you need 3 roots, 2 of which have to be complex. Usually complex roots are in pairs so assume -2i as well as 2i.

OpenStudy (anonymous):

ok i need u to show me not explain please i cant understand it if u just tell me

OpenStudy (anonymous):

(x-11)(x-2i)(x+2i)

OpenStudy (anonymous):

ok so i multiply everything together?

OpenStudy (anonymous):

Use the foil method to multiple the three factors. Multiply the first two factors first and then multiply that result by the third one. Just remember that i^2=-1

OpenStudy (anonymous):

thats the main thing i dont understand

OpenStudy (anonymous):

-2xi=(-2)i(x) this will subtract out when you multiply (x)(2i) Notice that (x-2i)(x+2i) is a difference of squares. So the product of these factors will be x^2-4i^2

OpenStudy (anonymous):

x^2- (11+2i)x + 22i is what i got

OpenStudy (anonymous):

There wouldn't be any i term. Note that (x-2i)(x+2i) =x^2+[(-4)(i^2) Which becomes x^2-4i^2 Since i^2=-1 It becomes x^2+4 Now multiply (x^2+4)(x-11)

OpenStudy (anonymous):

so the i just goes away?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

x^3-11x^2+4x-44

OpenStudy (anonymous):

Looks good to me.

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

One more thing. We are assuming the coefficients of the polynomial are real numbers. If you can have coefficients that are no real, then I believe you just multiply (x-2i)(x-11) x^2+(-2i-11)x+22i